### v2

# Supertranslation in the bulk for generic spacetime

Abstract
Supertranslations are usually defined as asymptotic symmetries associated with spacetime boundaries, such as null infinity and black hole horizons. In this Letter, we show that supertranslations admit a natural, coordinate-independent extension into the bulk of spacetime, realized as transitions between families of null hypersurfaces. This construction applies to generic spacetimes in arbitrary dimensions and unifies the realizations of supertranslations at null infinity and black hole horizons. The bulk supertranslation is connected to boundary supertranslation by characteristic flows. The associated symmetry algebra can be realized by light-ray operators defined on the null hypersurface. Within this framework, the gravitational wave memory effect corresponds to a shift of null hypersurfaces in the bulk. As explicit examples, we compute bulk supertranslations in Minkowski spacetime in arbitrary dimensions and in four-dimensional Schwarzschild spacetime, where we uncover a novel curvature-induced memory effect with observable consequences for light propagation.
Introduction. Supertranslation is a very surprising result from the investigation on asymptotic structure of the spacetime at null infinity Bondi et al. (1962); Sachs (1962a, b), where the spacetime symmetry is enhanced from the Poincaré group to the Bondi-Metzner-Sachs (BMS) group. The BMS group is the semidirect product of the Lorentz group and an infinite-dimensional Abelian group, the supertranslation that generalizes the ordinary translations. Although supertranslation symmetry was discovered long ago, its physical implications have only been appreciated recently Strominger (2018). The major breakthrough was made by Strominger for the discovery of the BMS invariance of gravitational scattering Strominger (2014). Supertranslations lie at the heart of a fascinating triangular equivalence Strominger (2018). It reveals the symmetry origin of the Weinberg’s soft graviton theorem at the quantum level He et al. (2015) and characterizes the vacua transition associated with the gravitational memory effect at the classical level Strominger and Zhiboedov (2016). Further developments from various perspectives can be found, e.g., in Carlip (2018); Bousso and Porrati (2017); Sousa et al. (2018); Choi and Akhoury (2018); Córdova and Shao (2018); Donnay et al. (2020); Chen et al. (2021); Fuentealba et al. (2021); Chakraborty et al. (2022); Fuentealba et al. (2022); Veneziano and Vilkovisky (2022); Javadinezhad and Porrati (2023); Fuentealba et al. (2023); Henneaux (2024); Javadinezhad and Porrati (2024); Elkhidir et al. (2025); De Luca et al. (2025a, b); Henneaux (2025).
A natural question is whether supertranslations or other asymptotic symmetries can arise in the bulk of spacetime Compère and Long (2016a, b); Compère (2016); CL . Important progress was achieved with the discovery of near-horizon symmetries Donnay et al. (2016a, b), but their emergence relies on inner boundaries such as black hole horizons, which share structural features with null infinityAshtekar and Speziale (2024a, b, c). Whether supertranslations can be defined in the bulk of generic spacetime has not been established. Resolving this question is essential, as bulk supertranslations could unify near-horizon and null-infinity symmetries and yield complete conservation laws, which are fundamental in the application of soft hairs in resolving the black hole information paradox Hawking (2015); Hawking et al. (2016) and in the description of vacua transition associated with the bulk gravitational memory Bart (2020). This raises the intriguing possibility that supertranslations are not merely asymptotic or near horizon symmetries but encode physical effects throughout the bulk of spacetime.
Supertranslations were originally derived with respect to particular gauge conditions in a specific coordinate system, such as the Bondi gauge Bondi et al. (1962); Sachs (1962a) or the Newman-Unti (NU) gauge Newman and Penrose (1962); Newman and Unti (1962) in the Newman-Penrose formalism Newman and Penrose (1962). Despite these differences, the gauges share the same asymptotic structure at null infinity, and supertranslations can be defined using purely geometric and conformal method Penrose (1963); Geroch (1977); Wald (1984); Ashtekar (2015) and references therein. Therefore, a proper definition of a bulk supertranslation requires not only specifying the physical structure it preserves but also formulating it in a coordinate-independent manner.
In this Letter, we demonstrate that bulk supertranslations can be defined as transitions between null hypersurfaces, generated by characteristic flows that uniquely extend boundary supertranslations into the spacetime interior. Since null hypersurfaces are geometric entities, this definition is inherently coordinate-independent. The bulk symmetry algebra can be realized by light-ray operators defined from light-ray integral of the energy-momentum tensor on the null hypersurface Córdova and Shao (2018). Those constructions can, in principle, be applied to any Lorentzian spacetime in arbitrary dimensions.
Motivated by the intrinsic connection between supertranslations and gravitational memory, we propose a bulk memory effect corresponding to the transition of null geodesics from one hypersurface to another. Thus, supertranslations are not merely asymptotic but define a genuine bulk symmetry with observable memory effects Goncharov et al. (2024). As concrete demonstrations, we compute supertranslations for Minkowski spacetime in four dimensions, which recovers the bulk extension of supertranslations in Minkowski spacetime obtained in Compère and Long (2016a). The computations are then extended to Minkowski spacetime in arbitrary dimensions.
A novel curvature-dependent memory effect is uncovered from the supertranslation in four-dimensional Schwarzschild spacetime, where elliptic integrals are involved with branch points that correspond to the turning points of null geodesics on the hypersurface. Specifically, light rays initially placed on a hypersurface without turning points can be shifted to a new hypersurface with turning points by gravitational wave memory, which allows them to return from the bulk. This type of memory effect occurs only in curved spacetime flat , namely black hole memory Donnay et al. (2018); Rahman and Wald (2020), and can be directly determined within black hole perturbation theory, such as gravitational self-force theory Elhashash and Nichols (2025); Cunningham et al. (2025).
The Newman-Unti coordinates. The NU gauge is based on a family of null hypersurface which can always be introduced in a normal hyperbolic Riemannian manifold Newman and Penrose (1962); Newman and Unti (1962). Suppose this family of null hypersurface is parametrized by u=u=const, and $u$ is chosen as one of the spacetime coordinates. The normal vector to those null hypersurface $\ell=\text{d}u$ must be null. Consequently, $\ell$ is tangent to null geodesics which lie within the null hypersurfaces.
The remarkable feature of foliating spacetime with null hypersurfaces is that one can formulate the Einstein equation or other perturbative equations on a background spacetime as a characteristic initial value problem Winicour (2001). More specifically, the equations of motion can be organized into two groups: evolution along $\ell$ on a single null hypersurface, and evolution between different hypersurfaces. It is very convenient to choose the geodesic parameter (e.g., $r$) as one coordinate of the spacetime, so that the evolution equations along a single null hypersurface reduce to ordinary differential equations, because the directional derivative along $\ell$ is proportional to $\frac{\partial}{\partial r}$. If $r$ is an affine parameter for the geodesic, $\ell=-\frac{\partial}{\partial r}$. The remaining two coordinates $x^{a}$ label the geodesics on each null hypersurface. This completes the construction of the celebrated Newman-Unti gauge with the metric assumptions,
$g^{uu}=0,\qquad g^{ur}=-1,\qquad g^{ua}=0.$ (1)
It is important to note that an affine parameter is not required to formulate the equations of motion as a characteristic initial value problem. In the seminal work of Bondi and collaborators Bondi et al. (1962), a luminosity parameter was employed where the metric component $g^{ur}$ is left free while the determinant of the metric $g^{ab}$ is fixed. This choice would simplify the evolution equations on the null hypersurface, see, e.g., Barnich and Lambert (2012); Chesler and Yaffe (2014); Geiller and Zwikel (2022, 2024).
Residual gauge transformation and supertranslation. In this section, we first examine the physical implications of residual gauge transformations within the NU gauge and identify the supertranslations in this context. We then generalize these definitions to a generic coordinate system. Suppose that two NU coordinate systems $(u,r,x^{a})$ and $(\bar{u},\bar{r},\bar{x}^{a})$ are related by $\bar{x}^{\mu}=\bar{x}^{\mu}(x^{\alpha})$. A key feature of the NU gauge is the choice of a null hypersurface labeled by a single coordinate. If the spacetime can be described in both NU coordinate systems, then the hypersurfaces u=u=const and $\bar{u}=$const are both null. Consequently, the full residual gauge transformations of the NU gauge are characterized by different families of null hypersurfaces hypersurface . In the NU coordinates $(u,r,x^{a})$, a residual gauge transformation is determined by a scalar potential $\bar{u}$, with the condition that $\bar{\ell}=\text{d}\bar{u}$ is a null vector.
The physical significance of a residual gauge transformation in the NU framework lies in the change of null hypersurfaces from u=u=const and $\bar{u}=$const. Consequently, the system can be formulated as a characteristic initial value problem based on different families of null hypersurfaces. The coordinate change itself is not fundamental, as it merely reflects a different parametrization of the null hypersurfaces. From a symmetry perspective, all residual gauge transformations preserve the formulation of the system as a characteristic initial value problem. Potential applications of this feature can arise in the context of null surface quantization Frolov (1978) and light-ray operators Kravchuk and Simmons-Duffin (2018).
Since $\bar{\ell}=\text{d}\bar{u}$ is a null vector, any rescaling $\Phi\ell$ remains null. This rescaling freedom can be used to write the null vector in the form $\bar{\ell}=\text{d}\bar{u}=\text{d}u+\text{d}f(r,x^{a})$, so that $\bar{u}=u+f(r,x^{a})$. Imposing the null condition then leads to the constraint for $f$ as
$g^{rr}(\partial_{r}f)^{2}-2\partial_{r}f+2g^{ra}\partial_{r}f\partial_{a}f+g^{ab}\partial_{a}f\partial_{b}f=0.$ (2)
This equation is precisely the condition $g^{\bar{u}\bar{u}}=0$ in the transformed NU coordinates. So any solution of (2) corresponds to a supertranslation arising from a finite BMS transformation Bondi et al. (1962); Sachs (1962a, b); Barnich and Troessaert (2016); Flanagan and Nichols (2024). In particular, the constraint (2) reduces to an ordinary differential equation for variable $r$ when considering a series expansion near the null infinity for asymptotically flat spacetime. The integration constant in $f$ at ${\cal O}(1)$ in a $1/r$ expansion reproduces the standard BMS supertranslation at null infinity Bondi et al. (1962); Sachs (1962a, b). All the subleading terms in $f$ are uniquely fixed by (2). Therefore, a supertranslation in the NU gauge throughout the bulk spacetime can be defined as any scalar function $f(r,x^{a})$ such that $\text{d}(u+f)$ remains null.
A few immediate remarks regarding the generic definition of supertranslation in the bulk are as follows:
  • We have introduced a rescaling $\Phi$ to specify a supertranslation. However, this rescaling has its own significance, as it relates to various generalizations of BMS symmetry, including superrotations and Weyl-BMS symmetries, see, e.g., Barnich and Troessaert (2010a, b, c, 2011, 2016); Compère et al. (2018); Campiglia and Laddha (2015); Campiglia and Peraza (2020); Flanagan and Nichols (2017, 2024); Freidel et al. (2021).
  • The NU gauge is usually formulated in four dimensions, but it can be generalized to arbitrary dimensions. The essential step is the selection of a family of null hypersurfaces. Consequently, the residual gauge transformations and supertranslations described above can be straightforwardly extended to higher dimensions.
  • The definition of supertranslation does not depend on a specific spacetime boundary. For instance, inserting solutions in a series expansion with a cosmological constant, the resulting supertranslations are consistent with the $\Lambda$-BMS transformations derived in Compère et al. (2019, 2020).
  • Applying a near-horizon series expansion, the supertranslation reduces to the near horizon supertranslations identified in Donnay et al. (2016a, b), and also Afshar et al. (2016); Shi and Mei (2017); Akhmedov and Godazgar (2017); Chandrasekaran et al. (2018); Haco et al. (2018); Grumiller et al. (2020a); Adami et al. (2020c, a); Chen et al. (2020); Adami et al. (2020b); Chandrasekaran et al. (2022); Adami et al. (2021) for other realizations. This generic definition of supertranslation thus unifies the null infinity and near-horizon symmetries, providing a natural framework for matching conserved quantities at different boundaries, see, e.g., relevant investigations in Hawking et al. (2016); Grumiller et al. (2020); Ruzziconi and Zwikel (2025a); Agrawal et al. (2025); Ruzziconi and Zwikel (2025b).
The definition of supertranslation above is given in NU coordinates. However, a specific coordinate system is not essential, since the key feature of the NU construction is simply the selection of a family of null hypersurfaces. The residual gauge transformations correspond precisely to the freedom in choosing these null hypersurfaces. Therefore, one can define residual gauge transformations with respect to the NU construction without reference to any particular coordinate system, and this definition naturally extends to spacetimes of arbitrary dimensions.
Definition 1: In a given coordinate system $x^{\mu}$, the residual gauge transformation is defined from any scalar function $H(x^{\mu})$ such that $\text{d}H(x^{\mu})$ is a null vector.
Supertranslations are defined as the set of all residual gauge transformations modulo rescalings. In asymptotically flat spacetimes, they can be specified using a normalized timelike coordinate $t$ at infinity, such that $(\frac{\partial}{\partial t})^{2}=-1$ there. A supertranslation can be defined as follows:
Definition 2: In a given coordinate system $(t,x^{i})$, a supertranslation is defined from any scalar function $h(x^{i})$ such that $\text{d}(t+h)$ is a null vector.
A Bondi-like coordinate $u=t+h$ can be introduced from the above definition so that $g^{uu}=0$. In this coordinate, the equations of motion can be formulated as a characteristic initial value problem. This definition uniquely specifies a supertranslation throughout the bulk of spacetime and is not restricted to the NU gauge. It can also be applied directly to solutions that are difficult to express in NU coordinates, such as the Kerr solution Fletcher and Lun (1996, 2003); Venter and Bishop (2006).
The scalar function that specifies a supertranslation satisfies a nonlinear, first-order partial differential equation (PDE), which can be solved using the method of characteristic flows Evans (2010); Levandosky (2002). In general, characteristic curves determine solutions only implicitly. However, the new variables defined along the characteristics can be treated as new coordinates. In this way, the full coordinate transformation corresponding to a supertranslation can be explicitly constructed. In the following sections, we will derive supertranslations using characteristic flows for several exact solutions.
Four-dimensional Minkowski with plane boundary. The Minkowski line-element with a plane boundary in flat null coordinates $(u,r,z,\bar{z})$ He and Mitra (2019) is given by
$\text{d}s^{2}=-2\text{d}u\text{d}r+2r^{2}\text{d}z\text{d}\bar{z}.$ (3)
The null condition for the scalar function then reads
$\partial_{r}f=\frac{1}{r^{2}}\partial_{z}f\partial_{\bar{z}}f.$ (4)
This is a first-order fully nonlinear PDE, which can be solved by the method of characteristics following closely Levandosky (2002). The equation can be organized as
$F(r,z,\bar{z},p_{r},p,\bar{p})=p_{r}-\frac{1}{r^{2}}p\bar{p}=0,$ (5)
where we introduce new variables $p_{r}=\partial_{r}f$, $p=\partial_{z}f$, and $\bar{p}=\partial_{\bar{z}}f$. The characteristic equations for this system are
$\begin{split}&\frac{\text{d}r}{\text{d}s}=1,\quad\frac{\text{d}z}{\text{d}s}=-\frac{\bar{p}}{r^{2}},\quad\frac{\text{d}\bar{z}}{\text{d}s}=-\frac{p}{r^{2}},\\ &\frac{\text{d}p}{\text{d}s}=0,\quad\frac{\text{d}\bar{p}}{\text{d}s}=0,\quad\frac{\text{d}f}{\text{d}s}=p\frac{\text{d}z}{\text{d}s}+\bar{p}\frac{\text{d}\bar{z}}{\text{d}s}+\partial_{r}f\frac{\text{d}r}{\text{d}s}.\end{split}$ (6)
Clearly, $p$ and $\bar{p}$ are constants along characteristics, which yields that
$z=\zeta+\frac{\bar{p}}{r},\qquad\bar{z}=\bar{\zeta}+\frac{p}{r},$ (7)
where $\zeta$ and $\bar{\zeta}$ are constants labeling the characteristic curves. Applying the differential equation (4), one can obtain that
$\frac{\text{d}f}{\text{d}s}=-\frac{p\bar{p}}{r^{2}},\qquad f=f_{0}(\zeta,\bar{\zeta})+\frac{p\bar{p}}{r}.$ (8)
Since $p$ and $\bar{p}$ are constants on the characteristic curves, they can be fixed by the initial data at the infinity $r\to\infty$. Hence, $p=\partial_{\zeta}f_{0}$ and $\bar{p}=\partial_{\bar{\zeta}}f_{0}$. Thus, we have solved the equation (4) implicitly.
The solution for a bulk supertranslation $f$ is determined from the boundary variables $(\zeta,\bar{\zeta})$ defined in (7). This corresponds precisely to the characteristic flow of a boundary supertranslation $f_{0}$. The null hypersurface characterized by $f$ is therefore the characteristic flow of a cross-section at null infinity Geroch (1977) specified by the boundary supertranslation $f_{0}$.
Considering $(\zeta,\bar{\zeta})$ as new coordinates, one can obtain the following relation of the two coordinate systems
$u=\bar{u}-f_{0}-\frac{\partial_{\zeta}f_{0}\partial_{\bar{\zeta}}f_{0}}{r},\qquad z=\zeta+\frac{\partial_{\bar{\zeta}}f_{0}}{r},$ (9)
which precisely reproduces the transformation derived in Compère and Long (2016a) for a supertranslated Minkowski spacetime. In Compère and Long (2016a), a radial shift $r=\bar{r}-\partial_{\zeta}\partial_{\bar{\zeta}}f_{0}$ was introduced, which corresponds to the integration constant in $\frac{\text{d}r}{\text{d}s}=1$. However, this shift does not affect the definition of the supertranslation.
We now pinpoint the supertranslation effect from the new coordinates $(\bar{u},\bar{r},\zeta,\bar{\zeta})$. Importantly, this is not merely the introduction of a new coordinate system, but rather the selection of a new family of null hypersurfaces with normal vector $\bar{\ell}=\text{d}\bar{u}$. In other words, the objects of interest change from null geodesics on the hypersurfaces uu=const to null geodesics on the hypersurfaces $\bar{u}$=const. Null geodesics lying on different hypersurfaces can exhibit distinct deviation properties, characterized by their shear. The two families of null hypersurfaces intersect null infinity at different cross-sections, which are related by a boundary supertranslation.
These features follow from a kinematical analysis of the local bulk geometry and its associated supertranslation. From a dynamical perspective, supertranslations are equivalent to gravitational memory Strominger and Zhiboedov (2016). In the bulk, gravitational waves with memory will distort null geodesics Bart (2020), effectively forcing their trajectories to transition from one null hypersurface to another as illustrated in Fig. 1.
Figure 1: The black wavy lines represent gravitational waves with memory. The blue straight lines denote null geodesics. Initially, the null geodesics are plane-fronted (green dashed line) with vanishing shear. As the gravitational waves pass, they deform this congruence, forcing the null geodesics onto a new null hypersurface. The individual geodesics remain straight, but the associated light front is no longer planar. When described using the original geodesic parameter $r$, the light front is sampled at different values of $r$ along different rays, which manifests itself as a non-vanishing shear of the null congruence.
As a closing remark for this section, we briefly discuss the symmetry algebra in the bulk. While there is no local energy–momentum tensor associated with the propagation of gravitational waves, their generation necessarily involves matter fields carrying a well-defined energy–momentum tensor. Gravitational wave memory can therefore be captured by nonlocal operators defined via light-ray integrals of the energy–momentum tensors along null hypersurfaces, namely light-ray operators. It has been shown that the algebra of such operators contains the BMS algebra Córdova and Shao (2018). This strongly suggests that the algebra of infinitesimal residual gauge transformations preserving the NU construction admits a realization in terms of light-ray operators on the null hypersurface.
Minkowski with sphere boundary in arbitrary dimensions. In four dimensions, the supertranslated Minkowski spacetime with a sphere boundary was obtained in Compère and Long (2016a) by applying a combined supertranslation and Weyl rescaling to the plane boundary Minkowski metric. In this section, we will apply the characteristics to obtain the supertranslation for Minkowski spacetime with a sphere boundary in arbitrary dimensions. The line-element is given by
$\text{d}s^{2}=-\text{d}u^{2}-2\text{d}u\text{d}r+r^{2}\gamma_{ab}(x^{c})\text{d}x^{a}\text{d}x^{b},$ (10)
where the Latin characters denote indices on the $d$-dimensional sphere. The null condition for the scalar function reads
$2\partial_{r}f=(\partial_{r}f)^{2}+\frac{1}{r^{2}}\gamma^{ab}\partial_{a}f\partial_{b}f,$ (11)
where $\partial_{a}=\frac{\partial}{\partial x^{a}}$. The above equation can be rewritten as
$F(r,x^{a},p_{r},p_{a})=p_{r}^{2}-2p_{r}+\frac{1}{r^{2}}(p_{a})^{2}=0,$ (12)
where we define $p_{r}=\partial_{r}f$ and $p_{a}=\partial_{a}f$. The characteristic equations for deriving the supertranslation are
$\begin{split}&\frac{\text{d}r}{\text{d}s}=2(p_{r}-1),\quad\frac{\text{d}{x}^{a}}{\text{d}s}=\frac{2}{r^{2}}p^{a},\quad\frac{\text{d}{p_{r}}}{\text{d}s}=\frac{2}{r^{3}}(p_{a})^{2},\\ &\frac{\text{d}p_{a}}{\text{d}s}=-\partial_{a}\gamma^{cd}p_{c}p_{d},\qquad\frac{\text{d}f}{\text{d}s}=\frac{\text{d}r}{\text{d}s}p_{r}+\frac{\text{d}{x}^{a}}{\text{d}s}p_{a}.\end{split}$ (13)
We can deduce the conservation of total angular momentum from the above characteristic equations, $\frac{\text{d}}{\text{d}s}(p_{a})^{2}=0$, which will significantly simplify the derivation of supertranslation. Let $(p_{a})^{2}=L^{2}$ on the characteristic curves. Then, the constraint equation gives
$p_{r}=1\pm\sqrt{1-\frac{L^{2}}{r^{2}}}.$ (14)
We will take the minus solution for the outgoing case with initial data given at infinity $r\to\infty$. The solution for the supertranslation is obtained as
$f=f_{0}({\bar{x}}^{a})+r\left(1-\sqrt{1-\frac{L^{2}}{r^{2}}}\right),$ (15)
where $L^{2}={\bar{\gamma}}^{ab}\partial_{{\bar{x}}^{a}}f_{0}\partial_{{\bar{x}}^{b}}f_{0}$. Here, $f_{0}$ represents the supertranslation at infinity, and ${\bar{x}}^{a}$ can be interpreted as new coordinates for the Minkowski spacetime with a supertranslation. The explicit coordinate transformation can be obtained with the precise form of the angular metric $\gamma_{ab}$.
Unlike Minkowski spacetime, where bulk supertranslations do not introduce qualitatively new structures, spacetime curvature leads to genuinely new features, which we now illustrate from the Schwarzschild solution.
Schwarzschild solution in four dimensions. The line-element of Schwarzschild solution is given by
$\text{d}s^{2}=-\left(1-\frac{2M}{r}\right)\text{d}u^{2}-2\text{d}u\text{d}r+r^{2}\gamma_{ab}(x^{c})\text{d}x^{a}\text{d}x^{b}.$ (16)
The null condition for the scalar function is
$2\partial_{r}f=\left(1-\frac{2M}{r}\right)(\partial_{r}f)^{2}+\frac{1}{r^{2}}\gamma^{ab}\partial_{a}f\partial_{b}f.$ (17)
Solving this equation via the method of characteristic flows gives
$f=f_{0}({\bar{x}}^{a})+\int_{+\infty}^{r}\frac{\text{d}r^{\prime}}{1-\frac{2M}{r^{\prime}}}\left(1-\frac{1}{\sqrt{1-\frac{L^{2}}{{r^{\prime}}^{2}}\left(1-\frac{2M}{r^{\prime}}\right)}}\right),$ (18)
where $L^{2}={\bar{\gamma}}^{ab}\partial_{{\bar{x}}^{a}}f_{0}\partial_{{\bar{x}}^{b}}f_{0}$. As before, $f_{0}$ represents the supertranslation at null infinity. This integral generally involves elliptic functions and can exhibit branch points depending on the supertranslation field $f_{0}$. These branch points are turning points of null geodesics on the associated null hypersurface,
$\frac{\text{d}r}{\text{d}s}=-2\sqrt{1-\frac{L^{2}}{{r}^{2}}\left(1-\frac{2M}{r}\right)}=0.$ (19)
Those branch points should correspond to points on the supertranslation horizon discovered in Compère and Long (2016b).
This feature highlights a remarkable property of supertranslations in curved spacetimes. Supertranslations in Minkowski spacetime can distort null rays without creating turning points. In curved spacetimes such as Schwarzschild, a supertranslation can map null geodesics from a hypersurface without turning points to a new hypersurface where turning points appear. This phenomenon gives rise to a novel gravitational memory effect.
One may initially select a family of null hypersurfaces on which the null geodesics exhibit no turning points. Gravitational waves with memory can then shift the system to a different null hypersurface where the light rays develop turning points. As a result, it becomes, in principle, possible to receive light rays returning from the bulk of spacetime. This constitutes a new type of gravitational memory effect that arises purely from spacetime curvature. Such memory effect can be directly computed within black hole perturbation theory, without the need to refer to a flat spacetime background or to null infinity.
A simple example of such a memory effect in Schwarzschild spacetime can be realized on the equatorial plane, $\theta=\frac{\pi}{2}$ equator . Initially, the null ray lies along uu=const with $L=0$. As the gravitational waves pass, they induce a nonzero $L$ as memory effect. At a later stage, the null ray follows a geodesic. Using spherical symmetry, one can choose new coordinates such that the final null ray remains on the equatorial plane, $\bar{\theta}=\frac{\pi}{2}$. Now we have $L\neq 0$, there is the famous photon sphere structure Synge (1966); Luminet (1979); Claudel et al. (2001), where the minimal turning point is located at the photon sphere radius $r_{\text{min}}=3M$. The entire process is illustrated in Fig. 2.
Figure 2: The black region represents the black hole. The black wavy lines indicate gravitational waves with memory, while the blue lines show the null geodesic before and after the memory effect. The memory effect can be observed from the propagation of one light ray.
Conclusions and perspectives. In this Letter, we propose a general framework for defining supertranslations in the bulk of spacetime. The framework is formulated in terms of null hypersurfaces and is therefore coordinate-independent, allowing it to be applied to generic spacetime. We present explicit expressions for bulk supertranslations in Minkowski spacetime and characterize the associated bulk memory effect that aligns with the triangle equivalence at null infinity Strominger (2018). In curved spacetimes, we uncover a novel memory effect in which supertranslations bend null geodesics, leading to observable changes in the propagation of light rays.
We determine the full residual gauge transformations associated with the choice of null hypersurfaces, which consist of supertranslations and a Weyl-type rescaling of null vectors. In the present work, we focus on the supertranslation sector. It is of independent interest to investigate the rescaling factor $\Phi$, which account for various extensions of the BMS symmetry. Another important direction is to clarify the relation between the bulk supertranslations constructed here and those arising in other contexts, such as scattering amplitudes Elkhidir et al. (2025) and holography Marjieh et al. (2022). A particular notable feature of our construction is that bulk supertranslations are realized as characteristic flows of boundary supertranslations. In the context of AdS/CFT, the gravity dual of a deformation by the composite operator $T\bar{T}$ Smirnov and Zamolodchikov (2017); Cavaglià et al. (2016) (also Jiang (2021); He et al. (2025) for reviews) corresponds to moving the boundary to a finite radial cutoff McGough et al. (2018); Kraus et al. (2018); Hartman et al. (2019); Guica and Monten (2021); Jafari et al. (2020); He et al. (2023, 2024); moving . The associated $T\bar{T}$ flow has been identified as a characteristic flow Hou (2023). These parallels suggest that our method of flowing asymptotic symmetries into the bulk is naturally compatible with the AdS/CFT perspective and merits further investigation.
Acknowledgments. The author would like to thank Hamed Adami, Jia Du, Shaoqi Hou, Delong Kong, Shahin Sheikh-Jabbari, Vahid Taghiloo, Yu Tian, and Hongbao Zhang for useful discussions. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 12475059 and No. 11935009, and by Tianjin University Self-Innovation Fund Extreme Basic Research Project Grant No. 2025XJ21-0007.